# Reducing Circuit Depth in Lindblad Simulation via Step-Size Extrapolation

### v2

@@ I Introduction
Our analysis is closest to that of [28, 20], but we target open-system (Lindblad) dynamics. In particular, [20, Lemma 4] establishes how the observable-estimation error scales with the Trotter step and, crucially, shows that Richardson or Chebyshev extrapolation can reduce circuit depth without worsening the dependence on the total simulation time $T$. We prove an analogous result for Lindblad evolutions via backward-error analysis. A key difference is that higher-order Suzuki–Trotter type formulas for closed-system simulation necessarily involve negative step sizes [5]; such negative intervals violate complete positivity in Lindblad dynamics. Consequently, we restrict attention to low-order splitting formulas, making our method a genuine extrapolation to $\tau\to 0$ rather than an interpolation across positive and negative steps.
@@
$\|\mathcal{L}\|\leq l,\quad l:=2\norm{H}+2\sum_{j=1}^{J}\norm{L_{j}}^{2}.$ (7)
@@ III First-order Approximation by a Kraus Form
One can verify that the scheme in (15) constitutes a first-order numerical method for simulating Lindblad dynamics [6]. That is, for sufficiently small step size $\tau$, the channel $\mathcal{K}$ approximates the exact evolution $e^{\tau\mathcal{L}}$ with an error that scales $l^{2}\tau^{2}$. More precisely, for a fixed initial state $\rho$, the local error satisfies
$\left\|\mathcal{K}[\rho]-e^{\tau\mathcal{L}}[\rho]\right\|=\mathcal{O}(l^{2}\tau^{2}),$ (16)
@@
$\norm{\mathcal{M}_{2}}\leq\left(\norm{H}+\frac{1}{2}\sum_{j}\norm{L_{j}}^{2}\right)^{2}=\Theta(l^{2}).$ (18)
@@
Definition 2 (Generating sequence for Kraus Form)
We define $c_{i,j,k}$ as non-negative real numbers such that $c_{0,0,k}=0\text{ for }k\geq 1$, $c_{i,j,0}=\delta_{j,0}l^{i}$ for $i,j\geq 0$, and the rest of the entries are generated from the recursion relations below:
@@
We then simulate $\hat{\rho}$ up to time $s=1$, with the operator norm of the Lindblad operator $\ell$ multiplied by $T$. Therefore, $\ell T$ will appear together in the simulation complexity. This scaling for a first-order method ($p=1$) is also consistent with that in [20, Sec. 2.3].
Corollary 5
@@
$T\geq\frac{\sqrt{2}\log(1/\varepsilon)\,\sqrt{\log\log(1/\varepsilon)}}{\,l\,\sqrt{\log l}\,}.$ (38)
@@ IV First-order Approximation by Hamiltonian Dilation
Theorem 11 (Local Error Expansion)
@@
$\norm{\rho^{(2k)}_{\text{R}}}\leq(J+1)\frac{l^{k}}{k!}.$ (46)
@@ V Numerical Experiments
Figure 2: Least-square based extrapolation of $\langle O\rangle(\tau)$ with data generated by first-order Kraus-form approximation. $N_{\mathrm{shots}}=2\times 10^{7}$. Left: Equidistant time steps. Right: Chebyshev time steps. The continuous curves represent degree-7 polynomials fit to 99 noisy data points, then extrapolated to $\tau=0$.
@@ VI Summary and further discussions
We have carried out a comprehensive bias–variance analysis of Richardson-style algorithmic error mitigation for first-order quantum algorithms that simulate Lindblad dynamics. Without invoking any a-priori smoothness assumptions, we proved that an $n=\Omega\bigl(\log(1/\varepsilon)\bigr)$-point extrapolator reduces the maximum circuit depth required to reach precision $\varepsilon$ from $\mathcal{O}((lT)^{2}/\varepsilon)$ to $\mathcal{O}\bigl((lT)^{2}\log^{2}(1/\varepsilon)\bigr)$, while preserving the standard $1/\varepsilon^{2}$ sampling complexity. Extensive numerical experiments confirmed that the predicted depth savings translate into measurable fidelity gains on noisy hardware. This framework opens the door to provably robust quantum simulations of dissipative dynamics on NISQ devices, an essential ingredient for many problems in quantum chemistry and physics.
@@ Appendix B Bounding the coefficients $c_{i,j,k}$
$c_{i,j,0}=\delta_{j,0}l^{i}\leq\frac{C^{i+0}_{1}\,C^{0}_{2}}{j!}=\frac{C^{i}_{1}}{j!}.$ (66)
@@
    $c_{0,j,k}$ $=\,\frac{B}{j}\cdot c_{0,j-1,k-1}+\delta_{j,1}\cdot\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{c_{p+1,j-1,k-p}}{j(p+1)!}\$(67)
    $\leq\frac{B}{j}\cdot\frac{C^{k-1}_{1}\,C^{k-1}_{2}}{(j-1)!}+\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{C_{1}^{k+1}\,C_{2}^{k-p}}{j!(p+1)!}$ (68)
    $\leq\frac{C_{1}^{k}\,C_{2}^{k}}{j!}\left(\frac{B}{C_{1}\,C_{2}}+\frac{l^{k+1}}{C^{k}_{1}\,C^{k}_{2}}+\sum_{p=1}^{k-j}\frac{C_{1}}{C^{p}_{2}\,(p+1)!}\right)$ (69)
@@
    We note that because of the assumption $l\geq 1$, we have $C_{1}\geq e+1$.To derive the last line notice that $\frac{B}{C_{1}}\leq 1$, and $\frac{1}{C_{2}}\leq\frac{1}{e+1}$. Similarly $C^{k}_{1}\,C^{k}_{2}\geq l^{2k}(e+1)^{3k}$, then $\frac{l^{k+1}}{C^{k}_{1}\,C^{k}_{2}}\leq\frac{1}{e+1}$. Also, $\sum_{p=1}^{k-j}\frac{C_{1}}{C_{2}^{p}(p+1!}\leq\sum_{p=1}^{\infty}\frac{C_{1}}{C_{2}^{p}(p+1)!}=C_{1}(e^{-C_{2}}-1)\leq\frac{C_{1}}{e^{C_{2}}}\leq\frac{1}{e+1}$.
@@
    $c_{i,j,k}=\,l\cdot c_{i-1,j,k}+B\cdot c_{i-1,j,k-1}+\delta_{j,0}\cdot\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{c_{i+p,j,k-p}}{(p+1)!}$ (71)
    $\leq l\cdot\frac{C_{1}^{i+k-1}\,C_{2}^{k}}{j!}+B\cdot\frac{C_{1}^{i+k-2}\,C_{2}^{k-1}}{j!}+\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{C_{1}^{i+k}\,C_{2}^{k-p}}{j!\,(p+1)!}$ (72)
    $\leq\frac{C_{1}^{i+k}\,C_{2}^{k}}{j!}\left(\frac{l}{C_{1}}+\frac{B}{C_{1}^{2}\,C_{2}}+\frac{l^{i+k}}{C_{1}^{i+k}\,C_{2}^{k}\,(k-j+1)!}+\sum_{p=1}^{k-j}\frac{1}{C^{p}_{2}\,(p+1)!}\right)$ (73)
@@
    $c_{i,k,k}$ $=l\cdot c_{i-1,k,k}$ (75)
    $\leq l\cdot\frac{C_{1}^{i+k-1}\,C_{2}^{k}}{k!}$ (76)
    $=\frac{C^{i+k}\,C_{2}^{k}}{k!}\cdot\frac{l}{C_{1}}\leq\frac{C^{i+k}\,C_{2}^{k}}{k!}.$ (77)
@@ Appendix C Bounding the derivatives of error expansions
$\|\Gamma_{k}(t)\|\leq\int_{0}^{t}B\ \|\Gamma_{k-1}(s)\|+\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k}\frac{\|\Gamma^{(p+1)}_{k-p}(s)\|}{(p+1)!}\,ds.$ (82)
@@
$\|\Gamma^{(i)}_{k}(t)\|\leq l\|\Gamma^{(i-1)}_{k}(t)\|+B\|\Gamma^{(i-1)}_{k-1}(t)\|+\frac{l^{k+i}}{(k+1)!}+\sum_{p=1}^{k}\frac{\|\Gamma^{(i+p)}_{k-p}(t)\|}{(p+1)!}.$ (84)
@@
$\bigl\|\Gamma_{k}^{(i)}(t)\bigr\|\leq P_{i,k}(t)=\sum_{j=0}^{k}c_{i,j,k}\,t^{j}\,,\quad c_{0,0,k}=0\text{ for }k\geq 1,\quad c_{i,j,0}=\delta_{j,0}l^{i}.$ (85)
@@
    $\|\Gamma_{k}(t)\|\leq\int_{0}^{t}\Bigl(B\sum_{j=0}^{k-1}c_{0,j,k-1}s^{j}+\frac{l^{k+1}}{(k+1)!}+\sum_{j=0}^{k-1}\frac{\sum_{p=1}^{\,k-j}c_{p+1,j,k-p}}{(p+1)!}s^{j}\Bigr)\,ds,$ (86)
@@
    $\sum_{j=0}^{k}c_{0,j,k}\ t^{j}=\sum_{j=0}^{k}\left(\frac{B}{j}\,c_{0,j-1,k-1}+\delta_{j,1}\,\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{c_{\,p+1,j-1,k-p}}{j\,(p+1)!}\right)\,t^{j},$ (87)
@@
    $\|\Gamma_{k}^{(i)}(t)\|\leq l\sum_{j=0}^{k}c_{i-1,j,k}t^{j}+B\sum_{j=0}^{k-1}c_{i-1,j,k-1}t^{j}+\frac{l^{k+i}}{(k+1)!}+\sum_{j=0}^{k-1}\frac{\sum_{p=1}^{k-j}c_{\,p+i,j,k-p}}{(p+1)!}t^{j},$ (88)
@@
    $\sum_{j=0}^{k}c_{i,j,k}\ t^{j}=lc_{i-1,k,k}\,t^{k}+\sum_{j=0}^{k-1}\left(l\,c_{i-1,j,k}+B\,c_{i-1,j,k-1}+\delta_{j,0}\,\frac{l^{k+i}}{(k+1)!}+\sum_{p=1}^{\,k-j}\frac{c_{p+i,j,k-p}}{(p+1)!}\right)t^{j}\,$, (89)
@@ Appendix D Error bounds
Proof.[Proof] For a fixed final time $t$, Lemma 1 provides the step-size expansion
@@ Appendix E The Lebesgue constant for the perturbed Chebyshev nodes
$\biggl|\frac{\widetilde{\gamma}_{j}}{\gamma_{j}}\biggr|\leq\Bigl(1+\frac{A_{P}}{n}\Bigr)\,(1+\delta_{\max})\,A\,n^{\frac{4}{\pi^{2}-4}}\leq C\,n^{\frac{4}{\pi^{2}-4}}\quad(n\geq 2).$ (110)
Since $\sum_{j=1}^{n+1}|\gamma_{j}|=\mathcal{O}(\log n)$ for Chebyshev nodes, we conclude $\sum_{j=1}^{n+1}|\widetilde{\gamma}_{j}|\leq C\,n^{\frac{4}{\pi^{2}-4}}\log n$.
@@ Appendix F Local error expansion of the dilated Hamiltonian approximation
Theorem 16 (Even–order bound)
@@
$\|\Gamma_{k}(t)\|\leq\int_{0}^{t}\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{l^{p+1}}{(p+1)!}\|\Gamma_{k-p}(s)\|+\frac{\|\Gamma^{(p+1)}_{k-p}(s)\|}{(p+1)!}\,ds.$ (135)
@@
$\|\Gamma^{(i)}_{k}(t)\|\leq l\|\Gamma^{(i-1)}_{k}(t)\|+\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{l^{p+1}}{(p+1)!}\|\Gamma^{(i-1)}_{k-p}(t)\|+\frac{\|\Gamma^{(i+p)}_{k-p}(t)\|}{(p+1)!}.$ (137)
@@
Definition 17 (Generating sequence for Hamiltonian dilation)
We define $c_{i,j,k}$ as non-negative real numbers such that $c_{0,0,k}=0\text{ for }k\geq 1$, $c_{i,j,0}=\delta_{j,0}l^{i}$ for $i,j\geq 0$, and the rest of the entries are generated from the recursion relations below:
$c_{0,j,k}=$ $\,\delta_{j,1}\cdot\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\cdot\frac{l^{p+1}}{j(p+1)!}\cdot c_{0,j-1,k-p}+\frac{c_{p+1,j-1,k-p}}{j(p+1)!},\,j=1,2,\cdots,k$ (138)
$c_{i,j,k}=$ $\,l\cdot c_{i-1,j,k}+\delta_{j,0}\cdot\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\frac{l^{p+1}}{(p+1)!}c_{i-1,j,k-p}+\frac{c_{i+p,j,k-p}}{(p+1)!},j=0,1,\cdots,k-1$ (139)
$c_{i,k,k}=$ $\,l\cdot c_{i-1,k,k}.$ (140)
Lemma 18
@@
    $c_{0,j,k}=$ $\delta_{j,1}\cdot\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\cdot\frac{l^{p+1}}{j(p+1)!}\cdot c_{0,j-1,k-p}+\frac{c_{p+1,j-1,k-p}}{j(p+1)!}$ (143)
    $\leq$ $\frac{l^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}l^{p}\frac{C_{1}\cdot C_{1}^{k-p}C_{2}^{k-p}}{j!(p+1)!}+\frac{C_{1}^{k+1}C_{2}^{k-p}}{j!(p+1)!}$ (144)
    $\leq$ $\frac{C_{1}^{k}C_{2}^{k}}{j!}(\frac{l^{k+1}}{C_{1}^{k}C_{2}^{k}}+\sum_{p=1}^{k-j}\frac{C_{1}}{C_{1}^{p}C_{2}^{p}(p+1)!}+\frac{C_{1}}{C_{2}^{p}(p+1)!})$ (145)
@@
    On the second line we used $l\geq J+1$, which is the condition we assumed for the Hamiltonian dilation approximation method for longer-time simulations. We derived the last line similar to how we did in Appendix B.
@@
    $c_{i,j,k}=$ $\,l\cdot c_{i-1,j,k}+\delta_{j,0}\cdot\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\frac{l^{p+1}}{(p+1)!}c_{i-1,j,k-p}+\frac{c_{i+p,j,k-p}}{(p+1)!},$ (147)
    $\leq$ $l\cdot\frac{C_{1}^{i+k-1}C_{2}^{k}}{j!}+\frac{l^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}l^{p}\frac{C_{1}\cdot C_{1}^{i+k-p-1}C_{2}^{k-p}}{j!(p+1)!}+\frac{C_{1}^{i+k}C_{2}^{k-p}}{j!(p+1)!}$ (148)
    $\leq$ $\frac{C_{1}^{i+k}C_{2}^{k}}{j!}\left(\frac{l}{C_{1}}+\frac{l^{i+k}}{C_{1}^{i+k}C_{2}^{k}}+\sum_{p=1}^{k-j}\frac{l^{p}}{C_{1}^{p}C_{2}^{p}(p+1)!}+\frac{1}{C_{2}^{p}(p+1)!}\right)$ (149)

### v3

@@ I Introduction
Our analysis is closest to that of [28, 36], but we target open-system (Lindblad) dynamics. In particular, [36, Lemma 4] establishes how the observable-estimation error scales with the Trotter step and, crucially, shows that Richardson or Chebyshev extrapolation can reduce circuit depth without worsening the dependence on the total simulation time $T$. We prove an analogous result for Lindblad evolutions via backward-error analysis. A key difference is that higher-order Suzuki–Trotter type formulas for closed-system simulation necessarily involve negative step sizes [5]; such negative intervals violate complete positivity in Lindblad dynamics. Consequently, we restrict attention to low-order splitting formulas, making our method a genuine extrapolation to $\tau\to 0$ rather than an interpolation across positive and negative steps.
@@
$\|\mathcal{L}\|\leq\ell,\quad\ell:=2\norm{H}+2\sum_{j=1}^{J}\norm{L_{j}}^{2}.$ (7)
@@ III First-order Approximation by a Kraus Form
One can verify that the scheme in (15) constitutes a first-order numerical method for simulating Lindblad dynamics [7]. That is, for sufficiently small step size $\tau$, the channel $\mathcal{K}$ approximates the exact evolution $e^{\tau\mathcal{L}}$ with an error that scales $\ell^{2}\tau^{2}$. More precisely, for a fixed initial state $\rho$, the local error satisfies
$\left\|\mathcal{K}[\rho]-e^{\tau\mathcal{L}}[\rho]\right\|=\mathcal{O}(\ell^{2}\tau^{2}),$ (16)
@@
$\norm{\mathcal{M}_{2}}\leq\left(\norm{H}+\frac{1}{2}\sum_{j}\norm{L_{j}}^{2}\right)^{2}=\Theta(\ell^{2}).$ (18)
@@
Definition 2 (Generating sequence for Kraus Form).
We define $c_{i,j,k}$ as non-negative real numbers such that $c_{0,0,k}=0\text{ for }k\geq 1$, $c_{i,j,0}=\delta_{j,0}\ell^{i}$ for $i,j\geq 0$, and the rest of the entries are generated from the recursion relations below:
@@
We then simulate $\hat{\rho}$ up to time $s=1$, with the operator norm of the Lindblad operator $\ell$ multiplied by $T$. Therefore, $\ell T$ will appear together in the simulation complexity. This scaling for a first-order method ($p=1$) is also consistent with that in [36, Sec. 2.3].
Corollary 5.
@@
$T\geq\frac{\sqrt{2}\log(1/\varepsilon)\,\sqrt{\log\log(1/\varepsilon)}}{\,\ell\,\sqrt{\log\ell}\,}.$ (38)
@@ IV First-order Approximation by Hamiltonian Dilation
Theorem 11 (Local Error Expansion).
@@
$\norm{\rho^{(2k)}_{\text{R}}}\leq(J+1)\frac{\ell^{k}}{k!}.$ (46)
@@ V Numerical Experiments
Figure 2: Least-squares based extrapolation of $\langle O\rangle(\tau)$ with data generated by first-order Kraus-form approximation. $N_{\mathrm{shots}}=2\times 10^{7}$. Left: Equidistant time steps. Right: Chebyshev time steps. The continuous curves represent degree-7 polynomials fit to 99 noisy data points, then extrapolated to $\tau=0$.
@@ VI Summary and further discussions
We have carried out a comprehensive bias–variance analysis of Richardson-style algorithmic error mitigation for first-order quantum algorithms that simulate Lindblad dynamics. Without invoking any a-priori smoothness assumptions, we proved that an $n=\Omega\bigl(\log(1/\varepsilon)\bigr)$-point extrapolator reduces the maximum circuit depth required to reach precision $\varepsilon$ from $\mathcal{O}((\ell T)^{2}/\varepsilon)$ to $\mathcal{O}\bigl((\ell T)^{2}\log^{2}(1/\varepsilon)\bigr)$, while preserving the standard $1/\varepsilon^{2}$ sampling complexity. Extensive numerical experiments confirmed that the predicted depth savings translate into measurable fidelity gains on noisy hardware. This framework opens the door to provably robust quantum simulations of dissipative dynamics on NISQ devices, an essential ingredient for many problems in quantum chemistry and physics.
@@ Appendix B Bounding the coefficients $c_{i,j,k}$
$c_{i,j,0}=\delta_{j,0}\ell^{i}\leq\frac{C^{i+0}_{1}\,C^{0}_{2}}{j!}=\frac{C^{i}_{1}}{j!}.$ (66)
@@
    $c_{0,j,k}$ $=\,\frac{B}{j}\cdot c_{0,j-1,k-1}+\delta_{j,1}\cdot\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{c_{p+1,j-1,k-p}}{j(p+1)!}\$(67)
    $\leq\frac{B}{j}\cdot\frac{C^{k-1}_{1}\,C^{k-1}_{2}}{(j-1)!}+\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{C_{1}^{k+1}\,C_{2}^{k-p}}{j!(p+1)!}$ (68)
    $\leq\frac{C_{1}^{k}\,C_{2}^{k}}{j!}\left(\frac{B}{C_{1}\,C_{2}}+\frac{\ell^{k+1}}{C^{k}_{1}\,C^{k}_{2}}+\sum_{p=1}^{k-j}\frac{C_{1}}{C^{p}_{2}\,(p+1)!}\right)$ (69)
@@
    We note that because of the assumption $\ell\geq 1$, we have $C_{1}\geq e+1$.To derive the last line notice that $\frac{B}{C_{1}}\leq 1$, and $\frac{1}{C_{2}}\leq\frac{1}{e+1}$. Similarly $C^{k}_{1}\,C^{k}_{2}\geq\ell^{2k}(e+1)^{3k}$, then $\frac{\ell^{k+1}}{C^{k}_{1}\,C^{k}_{2}}\leq\frac{1}{e+1}$. Also, $\sum_{p=1}^{k-j}\frac{C_{1}}{C_{2}^{p}(p+1!}\leq\sum_{p=1}^{\infty}\frac{C_{1}}{C_{2}^{p}(p+1)!}=C_{1}(e^{-C_{2}}-1)\leq\frac{C_{1}}{e^{C_{2}}}\leq\frac{1}{e+1}$.
@@
    $c_{i,j,k}=\,\ell\cdot c_{i-1,j,k}+B\cdot c_{i-1,j,k-1}+\delta_{j,0}\cdot\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{c_{i+p,j,k-p}}{(p+1)!}$ (71)
    $\leq\ell\cdot\frac{C_{1}^{i+k-1}\,C_{2}^{k}}{j!}+B\cdot\frac{C_{1}^{i+k-2}\,C_{2}^{k-1}}{j!}+\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{C_{1}^{i+k}\,C_{2}^{k-p}}{j!\,(p+1)!}$ (72)
    $\leq\frac{C_{1}^{i+k}\,C_{2}^{k}}{j!}\left(\frac{\ell}{C_{1}}+\frac{B}{C_{1}^{2}\,C_{2}}+\frac{\ell^{i+k}}{C_{1}^{i+k}\,C_{2}^{k}\,(k-j+1)!}+\sum_{p=1}^{k-j}\frac{1}{C^{p}_{2}\,(p+1)!}\right)$ (73)
@@
    $c_{i,k,k}$ $=\ell\cdot c_{i-1,k,k}$ (75)
    $\leq\ell\cdot\frac{C_{1}^{i+k-1}\,C_{2}^{k}}{k!}$ (76)
    $=\frac{C^{i+k}\,C_{2}^{k}}{k!}\cdot\frac{\ell}{C_{1}}\leq\frac{C^{i+k}\,C_{2}^{k}}{k!}.$ (77)
@@ Appendix C Bounding the derivatives of error expansions
$\bigl\|\Gamma_{k}^{(i)}(t)\bigr\|\leq P_{i,k}(t)=\sum_{j=0}^{k}c_{i,j,k}\,t^{j}\,\text{ for }i\geq 0,k\geq 1,\quad c_{0,0,k}=0,\quad c_{i,j,0}=\delta_{j,0}\ell^{i},$ (78)
@@
$\|\Gamma_{k}(t)\|\leq\int_{0}^{t}B\ \|\Gamma_{k-1}(s)\|+\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k}\frac{\|\Gamma^{(p+1)}_{k-p}(s)\|}{(p+1)!}\,ds.$ (82)
@@
$\|\Gamma^{(i)}_{k}(t)\|\leq\ell\|\Gamma^{(i-1)}_{k}(t)\|+B\|\Gamma^{(i-1)}_{k-1}(t)\|+\frac{\ell^{k+i}}{(k+1)!}+\sum_{p=1}^{k}\frac{\|\Gamma^{(i+p)}_{k-p}(t)\|}{(p+1)!}.$ (84)
@@
$\bigl\|\Gamma_{k}^{(i)}(t)\bigr\|\leq P_{i,k}(t)=\sum_{j=0}^{k}c_{i,j,k}\,t^{j}\,,\quad c_{0,0,k}=0\text{ for }k\geq 1,\quad c_{i,j,0}=\delta_{j,0}\ell^{i}.$ (85)
@@
    $\|\Gamma_{k}(t)\|\leq\int_{0}^{t}\Bigl(B\sum_{j=0}^{k-1}c_{0,j,k-1}s^{j}+\frac{\ell^{k+1}}{(k+1)!}+\sum_{j=0}^{k-1}\frac{\sum_{p=1}^{\,k-j}c_{p+1,j,k-p}}{(p+1)!}s^{j}\Bigr)\,ds,$ (86)
@@
    $\sum_{j=0}^{k}c_{0,j,k}\ t^{j}=\sum_{j=0}^{k}\left(\frac{B}{j}\,c_{0,j-1,k-1}+\delta_{j,1}\,\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\frac{c_{\,p+1,j-1,k-p}}{j\,(p+1)!}\right)\,t^{j},$ (87)
@@
    $\|\Gamma_{k}^{(i)}(t)\|\leq\ell\sum_{j=0}^{k}c_{i-1,j,k}t^{j}+B\sum_{j=0}^{k-1}c_{i-1,j,k-1}t^{j}+\frac{\ell^{k+i}}{(k+1)!}+\sum_{j=0}^{k-1}\frac{\sum_{p=1}^{k-j}c_{\,p+i,j,k-p}}{(p+1)!}t^{j},$ (88)
@@
    $\sum_{j=0}^{k}c_{i,j,k}\ t^{j}=\ell c_{i-1,k,k}\,t^{k}+\sum_{j=0}^{k-1}\left(\ell\,c_{i-1,j,k}+B\,c_{i-1,j,k-1}+\delta_{j,0}\,\frac{\ell^{k+i}}{(k+1)!}+\sum_{p=1}^{\,k-j}\frac{c_{p+i,j,k-p}}{(p+1)!}\right)t^{j}\,$, (89)
@@ Appendix D Error bounds
Proof. For a fixed final time $t$, Lemma 1 provides the step-size expansion
@@ Appendix E The Lebesgue constant for the perturbed Chebyshev nodes
$\biggl|\frac{\widetilde{\gamma}_{j}}{\gamma_{j}}\biggr|\leq\Bigl(1+\frac{A_{P}}{n}\Bigr)\,(1+\delta_{\max})\,A\,n^{4/(\pi^{2}-4)}\leq C\,n^{4/(\pi^{2}-4)}\quad(n\geq 2).$ (110)
Since $\sum_{j=1}^{n+1}|\gamma_{j}|=\mathcal{O}(\log n)$ for Chebyshev nodes, we conclude $\sum_{j=1}^{n+1}|\widetilde{\gamma}_{j}|\leq C\,n^{4/(\pi^{2}-4)}\log n$.
@@ Appendix F Local error expansion of the dilated Hamiltonian approximation
Theorem 16 (Even–order bound).
@@
$\|\Gamma_{k}(t)\|\leq\int_{0}^{t}\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{\ell^{p+1}}{(p+1)!}\|\Gamma_{k-p}(s)\|+\frac{\|\Gamma^{(p+1)}_{k-p}(s)\|}{(p+1)!}\,ds.$ (135)
@@
$\|\Gamma^{(i)}_{k}(t)\|\leq\ell\|\Gamma^{(i-1)}_{k}(t)\|+\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k}(J+1)\frac{\ell^{p+1}}{(p+1)!}\|\Gamma^{(i-1)}_{k-p}(t)\|+\frac{\|\Gamma^{(i+p)}_{k-p}(t)\|}{(p+1)!}.$ (137)
@@
Definition 17 (Generating sequence for Hamiltonian dilation).
We define $c_{i,j,k}$ as non-negative real numbers such that $c_{0,0,k}=0\text{ for }k\geq 1$, $c_{i,j,0}=\delta_{j,0}\ell^{i}$ for $i,j\geq 0$, and the rest of the entries are generated from the recursion relations below:
$c_{0,j,k}=$ $\,\delta_{j,1}\cdot\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\cdot\frac{\ell^{p+1}}{j(p+1)!}\cdot c_{0,j-1,k-p}+\frac{c_{p+1,j-1,k-p}}{j(p+1)!},\,j=1,2,\cdots,k$ (138)
$c_{i,j,k}=$ $\,\ell\cdot c_{i-1,j,k}+\delta_{j,0}\cdot\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\frac{\ell^{p+1}}{(p+1)!}c_{i-1,j,k-p}+\frac{c_{i+p,j,k-p}}{(p+1)!},j=0,1,\cdots,k-1$ (139)
$c_{i,k,k}=$ $\,\ell\cdot c_{i-1,k,k}.$ (140)
Lemma 18.
@@
    $c_{0,j,k}=$ $\delta_{j,1}\cdot\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\cdot\frac{\ell^{p+1}}{j(p+1)!}\cdot c_{0,j-1,k-p}+\frac{c_{p+1,j-1,k-p}}{j(p+1)!}$ (143)
    $\leq$ $\frac{\ell^{k+1}}{(k+1)!}+\sum_{p=1}^{k-j}\ell^{p}\frac{C_{1}\cdot C_{1}^{k-p}C_{2}^{k-p}}{j!(p+1)!}+\frac{C_{1}^{k+1}C_{2}^{k-p}}{j!(p+1)!}$ (144)
    $\leq$ $\frac{C_{1}^{k}C_{2}^{k}}{j!}(\frac{\ell^{k+1}}{C_{1}^{k}C_{2}^{k}}+\sum_{p=1}^{k-j}\frac{C_{1}}{C_{1}^{p}C_{2}^{p}(p+1)!}+\frac{C_{1}}{C_{2}^{p}(p+1)!})$ (145)
@@
    On the second line we used $\ell\geq J+1$, which is the condition we assumed for the Hamiltonian dilation approximation method for longer-time simulations. We derived the last line similar to how we did in Appendix B.
@@
    $c_{i,j,k}=$ $\,\ell\cdot c_{i-1,j,k}+\delta_{j,0}\cdot\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}(J+1)\frac{\ell^{p+1}}{(p+1)!}c_{i-1,j,k-p}+\frac{c_{i+p,j,k-p}}{(p+1)!},$ (147)
    $\leq$ $\ell\cdot\frac{C_{1}^{i+k-1}C_{2}^{k}}{j!}+\frac{\ell^{i+k}}{(k+1)!}+\sum_{p=1}^{k-j}\ell^{p}\frac{C_{1}\cdot C_{1}^{i+k-p-1}C_{2}^{k-p}}{j!(p+1)!}+\frac{C_{1}^{i+k}C_{2}^{k-p}}{j!(p+1)!}$ (148)
    $\leq$ $\frac{C_{1}^{i+k}C_{2}^{k}}{j!}\left(\frac{l}{C_{1}}+\frac{\ell^{i+k}}{C_{1}^{i+k}C_{2}^{k}}+\sum_{p=1}^{k-j}\frac{\ell^{p}}{C_{1}^{p}C_{2}^{p}(p+1)!}+\frac{1}{C_{2}^{p}(p+1)!}\right)$ (149)